OpenStudy (anonymous):

At 3 pm, a boat is 12.5 miles due west of radar station and traveling at 11 mph in a direction that's 57.3 degrees south of an east-west line. At what time will the boat be closest to the radar station?

OpenStudy (anonymous):

Please help me with this question

OpenStudy (aum):



OpenStudy (aum):

Find x. Divide x by the speed of the boat to find the time it will take for the boat to reach point C which will be the closest point to the radar station. Add this time to 3:00 pm to get the time the boat will be closest to the radar station.

OpenStudy (aum):

What trigonometric function connects side BC and side AB of the triangle ABC?

OpenStudy (aum):

cos(57.3) = adjacent / hypotenuse = BC / AB = x / 12.5 x = 12.5 * cos(57.3) = ?

OpenStudy (anonymous):

Yes, cosine will be used

OpenStudy (radar):

@ice0Learner Welcome to open study. Continue with the problem, you have been provided excellent assistance. Try and solve the problem.

OpenStudy (anonymous):

Thanks :)